Show Posts

This section allows you to view all posts made by this member. Note that you can only see posts made in areas you currently have access to.


Topics - misterpants

Pages: [1]
1
Hi folks--

I'm building a Mudbunny board with the Creamy Dreamer BOM. When the tone control is engaged, it works beautifully and sounds great. When I use the tone bypass, however, I lose all my signal.

I've tried changing the settings on the dip switches, but I get the same result. I verified the switch with a continuity tester, and everything looks good.

Has anyone else encountered this issue? If so, what steps did you take to resolve the problem?

2
Build Reports / Pepper Spray with mystery meat controls
« on: September 18, 2020, 11:32:18 AM »
Here's a Pepper Spray I just completed for a friend. He asked for mystery controls, so I just brushed a plain 1590B for the enclosure.

It's built to the 'Albini' spec. For the second set of clipping diodes, I used one 1N914 and one BAT46. I put a small 15N ceramic cap around the BAT46 to turn the second diode set into a switchable low pass filter. The BAT46 helps keep the volume fairly consistent when switching between diode sets.

The 404A is a nice-sounding old Texas Instruments transistor with an HFE of about 140. The 3565 has an HFE of about 260.

It's a surprising pedal, much warmer and subtler than the Percolators I've built to the 'stock' specs.

3
Tech Help - Projects Page / Pepper Spray Troubleshooting
« on: July 26, 2020, 12:59:54 PM »
I am building the 2019 version of the Pepper Spray, and I'm presently getting almost no output from the pedal. All I get are short, quiet staccato bursts of very distorted signal.

There are no part substitutions in the build. I am using socketed 1N914 diodes for D3 and D4. Q1 is a 2n404a from Small Bear, and Q2 is a 2N3565 from Small Bear.

Here is what I have done thus far:

  • Verified the orientation of all diodes
  • Verified the orientation and values of all polarized caps
  • Verified the values of all non-polarized caps
  • Verified the values of all resistors
  • Used an audio probe to trace the signal
  • Measured the DC voltages on Q1 and Q2 with a multimeter

The audio probe found a viable signal at the 100K harmonic pot, but almost nothing at the 50K balance pot.

The voltages are screwy.

Q1-C: 3.0mV
Q1-B: 16.2mV
Q1-E: 7.41v

Q2-C: 9.16v
Q2-B: 7.3v
Q2-E: 7.3v

I've attached a picture of the top of the board. It's hard to tell from the photo, but the flat part of Q2 (2N3565) is pointing toward the trim pot.

Any suggestions for next steps in debugging this are welcome.

Pages: [1]